# Existence of transmission eigenvalues for biharmonic scattering by a clamped planar region

### v1

@@ Abstract
In this paper, we study the so-called clamped transmission eigenvalue problem. This is a new transmission eigenvalue problem that is derived from the scattering of an impenetrable clamped obstacle in a thin elastic plate. The scattering problem is modeled by a biharmonic wave operator given by the Kirchhoff–Love infinite plate problem in the frequency domain. These scattering problems have not been studied to the extent of other models. Unlike other transmission eigenvalue problems, the problem studied here is a system of homogeneous PDEs defined in all of $\mathbb{R}^{2}$. This provides unique analytical and computational difficulties when studying the clamped transmission eigenvalue problem. We are able to prove that there exist infinitely many real clamped transmission eigenvalues. This is done by studying the equivalent variational formulation. We also investigate the relationship of the clamped transmission eigenvalues to the Dirichlet and Neumann eigenvalues of the negative Laplacian for the bounded scattering obstacle.
@@ 1 Introduction
In this paper, we provide an analytical and numerical study of the transmission eigenvalues associated with the scattering of an impenetrable clamped obstacle in a thin elastic plate. The scattering problem is modeled by the Kirchhoff–Love infinite plate equation, which leads to a biharmonic scattering problem. Recently, in [5, 16] the authors studied the well–posedness for the direct scattering in a thin elastic plate for a clamped obstacle. Here, we study a new transmission eigenvalue problem associated with the aforementioned scattering problem. The transmission eigenvalue problem studied here is different than previously studied problems. This is due to the fact that, this eigenvalue problem is posed on all of $\mathbb{R}^{2}$ despite the scatterer being a bounded domain. This makes the study of these eigenvalues challenging. For instance, since the problem is posed in an unbounded domain we no longer have the compact embedding of the standard Sobolev spaces. Transmission eigenvalue problems are often non–self–adjoint and nonlinear, which adds to the difficulty in studying them analytically and numerically. The main contribution of this work is proving the existence of the so–called clamped transmission eigenvalues as well as developing new theoretical techniques that can be used for exterior transmission eigenvalue problem [8, 13, 27].
@@
As in previous works, we see that this eigenvalue problem is derived from studying the associated inverse scattering problem. While inverse scattering problems have been extensively studied in acoustic and electromagnetic scattering theory, the corresponding problems for biharmonic scattering has only recently begun to attract attention [6, 12, 17, 19, 21, 28, 30]. The aforementioned manuscripts have studied the inverse shape problem given either near or far field data for a few kinds of scatterers. In the case of using far field data, it has been shown that the far field operator associated with a clamped obstacle fails to be injective with a dense range at a transmission eigenvalues see for e.g. [20]. Although these inverse problems are relevant in various applications, including non–destructive testing and medical imaging [14, 15], both theoretical and numerical investigations remain relatively limited.
@@
Transmission eigenvalue problems have played an important role in inverse scattering theory. The existence of interior transmission eigenvalues for an isotropic acoustic scatterer has been studied in [3, 9, 10], and the exterior transmission eigenvalues for an isotropic acoustic scatterer has been investigated in [13] for a spherically stratified media in three dimensions. Throughout the years, the associated inverse spectral problem has been of interest. This is due to the fact that these eigenvalues can be shown to depend (often monotonically) on the material properties of the scatterer. This has given rise to studying the inverse spectral problem to recover information about the scatterer from the transmission eigenvalues. In general, it has been shown that for acoustic and electromagnetic scatterers that these eigenvalues depend monotonically on the material parameters and geometric properties of the scatterer, see for e.g. [3, 11, 23].
@@
This paper is organized as follows. In Section 2, we formulate the direct scattering problem and associated clamped transmission eigenvalue problem. These eigenvalues can be seen as the frequencies where there is an incident wave that produces trivial far field data. In Section 3, we prove the existence of real transmission eigenvalues using the techniques introduced in [10]. Additionally, we study the relationship between the first transmission eigenvalue and the first Dirichlet eigenvalue of the Laplacian. We then numerically investigate the interlacing between the clamped transmission eigenvalues along with the Dirichlet and Neumann eigenvalues for the obstacle in Section 4. Lastly, Section 5 provides a summary of our work along with outlooks to other interesting problems in this direction.
@@ 2 Formulation of Transmission Eigenvalue Problem
In this section, we introduce the biharmonic scattering problem associated with the clamped transmission eigenvalues. Here, we assume that the clamped obstacle is modeled by a bounded region $D\subset\mathbb{R}^{2}$ with an analytic boundary $\partial D$. To illuminate the obstacle, one uses a time–harmonic incident plane wave, which we denote as $u^{\text{inc}}(x)=\text{e}^{\text{i}kx\cdot d}$, where $d$ is the corresponding incident direction such that $d\in\mathbb{S}^{1}=\{x\in\mathbb{R}^{2}\mathrel{\mathop{\ordinarycolon}}|x|=1\}$. For our model, the corresponding scattered field $u^{\text{scat}}$ satisfies the biharmonic scattering problem for a fixed wave number $k>0$, given by
@@ 3 Existence of the Transmission Eigenvalues
We note that, for $k>0$ we have that both $u_{f}$ and $\partial_{r}u_{f}$ decay exponentially as $r\to\infty$ so there is no radiation condition at infinity but we assume that $u_{f}\in H^{1}(\mathbb{R}^{2}\setminus\overline{B_{R}})$. First, note that the eigenvalue problems (7)–(8) and (9)–(10) are equivalent.
@@
This implies that we can prove the the existence of eigenpairs $k>0$ and $(v,w)\in X(D,B_{R})$ to (9)–(10) where the Hilbert space
@@
with the associated inner–product
@@
Therefore, to study the transmission eigenvalue problem in the truncated domain, we will consider the equivalent variational formulation of (9)–(10). By appealing to Green’s first identity we have that the equivalent variational formulation of the clamped transmission eigenvalue problem (9)–(10) is given by: find eigenpairs $k>0$ and $(v,w)\in X(D,B_{R})$ such that
@@
Notice, that $k>0$ is a clamped transmission eigenvalue if and only if the operator $\mathbb{A}_{k}-k^{2}\mathbb{B}$ has a non–trivial null space.
As previously stated, in [20] it has been shown that the set of clamped transmission eigenvalues $k>0$ corresponding to (7)–(8) (or equivalently (9)–(10)) is at most a discrete. With this, we focus on the existence of clamped transmission eigenvalues in this work. To prove this claim, we will study the analytic properties of the operators $\mathbb{A}_{k}$ and $\mathbb{B}$ defined in (16)–(17). Motivated by the results in [10], we will show that the aforementioned operators satisfy the assumptions of a key result. To this end, we recall Theorem 2.3 of [10] which will be applied to our eigenvalue problem.
@@
This result is proven by considering the problem $\lambda_{j}(k)-k^{2}=0$ where $\lambda_{j}(k)$ are the generalized eigenvalues such that $\mathbb{A}_{k}-\lambda_{j}(k)\mathbb{B}$ has a non–trivial null space. Since each $\lambda_{j}(k)$ is continuous with respect to $k\geq 0$, the positivity and non–positivity requirements in Lemma 3.1 implies that
@@
Notice, that $C_{k}$ can depend on $k>0$ but not $f$ and we make the dependance of the solution with respect to $k$ explicit. We will now employ a variational argument to prove that
@@
Now, to continue we let $u^{k}_{f}$ and $u^{\tau}_{f}$ satisfy the above boundary value problem with ‘wave number’ $k>0$ and $\tau>0$, respectively. Notice that
@@
Therefore, we can now estimate to obtain
@@
With the above analysis we can now prove that the operators $\mathbb{A}_{k}$ and $\mathbb{B}$ satisfies the assumptions of Lemma 3.1. We prove that each operator satisfies the assumptions in the following two results.
@@
$\left|\left(\left(\mathbb{A}_{{k}}-\mathbb{A}_{{\tau}}\right)(v,w);(\psi_{1},\psi_{2})\right)_{X(D,B_{R})}\right|=\left|\int_{B_{R}\setminus\overline{D}}\left(k^{2}-\tau^{2}\right)w\overline{\psi_{2}}\,\mathrm{d}x+\int_{\partial B_{R}}\overline{\psi_{2}}\big{(}T_{\text{i}k}-T_{\text{i}\tau}\big{)}w\,\mathrm{d}s\right|.$
@@
where $\|\cdot\|$ is the appropriate operator norms. By Theorem 3.2 we have that $\|\mathbb{A}_{{k}}-\mathbb{A}_{{\tau}}\|\longrightarrow 0$ as $\tau\to k$, proving the continuity.
Now, to prove that the operator is self–adjoint. By appealing to the definition of the sesquilinear form in (14) we notice that
$\left(\mathbb{A}_{k}(v,w);(v,w)\right)_{X(D,B_{R})}$ $=\int_{D}|\nabla v|^{2}\,\mathrm{d}x+\int_{B_{R}\setminus\overline{D}}{|\nabla w|^{2}+k^{2}|w|}\,\mathrm{d}x-\int_{\partial B_{R}}\overline{w}T_{\text{i}k}w\,\mathrm{d}s$
$=\int_{D}|\nabla v|^{2}\,\mathrm{d}x+\int_{B_{R}\setminus\overline{D}}{|\nabla w|^{2}+k^{2}|w|}\,\mathrm{d}x+\int_{\mathbb{R}^{2}\setminus\overline{D}}{|\nabla u_{w}|^{2}+k^{2}|u_{w}|^{2}}\mathrm{d}x.$
@@
Since the sequence is bounded, we have that (up to a subsequence) it is weakly convergent to some $(v,w)\in{X(D,B_{R})}$. This implies that $w_{n}\to w=0$ in $H^{1}(B_{R}\setminus\overline{D})$ and that $|\nabla v_{n}|\to|\nabla v|=0$ in $L^{2}(D)$ as $n\to\infty$. Therefore, we see that the limiting value $v$ is constant. By appealing to the compact embedding $H^{1}(D)$ into $L^{2}(D)$ we have that (up to a subsequence) $v_{n}$ is strongly convergent in $H^{1}(D)$. The boundary condition v=wv=-w on $\partial D$ implies that $v=0$ in $D$. The convergence $(v_{n},w_{n})\to(0,0)$ in $X(D,B_{R})$ as $n\to\infty$ contradicts the fact that $\|(v_{n},w_{n})\|_{X(D,B_{R})}=1$. This proves the coercivity and therefore proves the claim. ∎
@@
It is clear that from the sequilinear form (15) that $\mathbb{B}$ given by (17) is self–adjoint and non–negative since $\mathcal{B}\big{(}(v,w)\,;(v,w)\big{)}=\|v\|^{2}_{L^{2}(D)}$. For the compactness, by definition we can easily obtain that $\|\mathbb{B}(v,w)\|_{X(D,B_{R})}\leq\|v\|_{L^{2}(D)}$. The compact embedding of $H^{1}(D)$ into $L^{2}(D)$ proves the claim, since $D$ is a bounded domain with an analytic boundary. ∎
@@
where $H^{(1)}_{\ell}$ denotes the first kind Hankel function of order $\ell$ and $f_{n}$ are the Fourier coefficients for $f\in H^{1/2}(\partial B_{R})$. Note that above we have used the fact that
@@ 4 Numerical Investigation
is the radiating fundamental solution for the Helmholtz equation in $\mathbb{R}^{2}$. We that $\lambda$ and $\mu$ are the corresponding Dirichlet and Neumann eigenvalues when
@@ 5 Conclusion
In conclusion, we have provided an analytical and numerical study of the new clamped transmission eigenvalue problem. This eigenvalue problem is associated with scattering in a thin elastic plate. Even though the scatterer is bounded, the eigenvalue problem (7)–(8) is posed in the entirety of $\mathbb{R}^{2}$. With the results in this paper, we now know that there exists an infinite discrete set of real eigenvalues such that $k>0$. Here, we have provided more justification that the first clamped transmission eigenvalue is monotonically decreasing with respect to the measure of $D$, but it has not yet been proved. Another question that arises from Theorem 3.8 is if there are interlacing inequalities with other eigenvalues associated with the scatterer, which we have numerically investigated in Section 4. Also, the eigenvalue problem studied here is derived via a clamped obstacle assumption. It would be interesting to see what eigenvalue problems one derives from either the simply supported or free plate obstacle assumption.

### v2

@@ Abstract
In this paper, we study the so-called clamped transmission eigenvalue problem. This is a new transmission eigenvalue problem that is derived from the scattering of an impenetrable clamped obstacle in a thin elastic plate. The scattering problem is modeled by a biharmonic wave operator given by the Kirchhoff–Love infinite plate problem in the frequency domain. These scattering problems have not been studied to the extent of other models. Unlike other transmission eigenvalue problems, the problem studied here is a system of homogeneous PDEs defined in all of $\mathbb{R}^{2}$. This provides unique analytical and computational difficulties when studying the clamped transmission eigenvalue problem. We are able to prove that there exist infinitely many real clamped transmission eigenvalues. This is done by studying the equivalent variational formulation. We also investigate the relationship of the clamped transmission eigenvalues to the Dirichlet and Neumann eigenvalues of the negative Laplacian for the bounded scattering obstacle.ă
@@ 1 Introduction
In this paper, we provide an analytical and numerical study of the transmission eigenvalues associated with the scattering of an impenetrable clamped obstacle in a thin elastic plate. The scattering problem is modeled by the Kirchhoff–Love infinite plate equation, which leads to a biharmonic scattering problem. Recently, in [6, 20] the authors studied the well–posedness for the direct scattering in a thin elastic plate for a clamped obstacle. Here, we study a new transmission eigenvalue problem associated with the aforementioned scattering problem that is different than previously studied problems [4, 9, 10, 27]. This is due to the fact that, this eigenvalue problem is posed on all of $\mathbb{R}^{2}$ despite the scatterer being a bounded domain. This makes the study of these eigenvalues challenging. For instance, since the problem is posed in an unbounded domain we no longer have the compact embedding of the standard Sobolev spaces. Transmission eigenvalue problems are often non–self–adjoint and nonlinear, which adds to the difficulty in studying them analytically and numerically. The main contribution of this work is proving the existence of the so–called clamped transmission eigenvalues as well as developing new theoretical techniques that can be used for exterior transmission eigenvalue problems [9, 16, 31].
@@
As in previous works [4, 9, 10, 27], we see that this eigenvalue problem is derived from studying the associated inverse scattering problem. While inverse scattering problems have been extensively studied in acoustic and electromagnetic scattering theory, the corresponding problems for biharmonic scattering has only recently begun to attract attention [7, 13, 21, 23, 25, 32, 34]. The aforementioned manuscripts have studied the inverse shape problem given either near or far field data for a few kinds of scatterers. In the case of using far field data, it has been shown that the far field operator associated with a clamped obstacle fails to be injective with a dense range at a transmission eigenvalues; see, e.g., [24]. Although these inverse problems are relevant in various applications, including non–destructive testing and medical imaging [17, 18], both theoretical and numerical investigations remain relatively limited.
@@
Transmission eigenvalue problems have played an important role in inverse scattering theory. The existence of interior transmission eigenvalues for an isotropic acoustic scatterer has been studied in [4, 10, 11], and the exterior transmission eigenvalues for an isotropic acoustic scatterer has been investigated in [16] for a spherically stratified media in three dimensions. Throughout the years, the associated inverse spectral problem has been of interest. This is due to the fact that these eigenvalues can be shown to depend (often monotonically) on the material properties of the scatterer. This has given rise to studying the inverse spectral problem to recover information about the scatterer from the transmission eigenvalues. In general, it has been shown that for acoustic and electromagnetic scatterers that these eigenvalues depend monotonically on the material parameters and geometric properties of the scatterer; see, e.g., [4, 12, 27].
@@
This paper is organized as follows. In Section 2, we formulate the direct scattering problem and associated clamped transmission eigenvalue problem. These eigenvalues can be seen as the frequencies where there is an incident wave that produces trivial far field data. In Section 3, we prove the existence of real transmission eigenvalues using the techniques introduced in [11]. Additionally, we study the relationship between the first transmission eigenvalue and the first Dirichlet eigenvalue of the negative Laplacian. We then numerically investigate the interlacing between the clamped transmission eigenvalues along with the Dirichlet and Neumann eigenvalues for the obstacle in Section 4. Lastly, Section 5 provides a summary of our work along with outlooks to other interesting problems in this direction.
@@ 2 Formulation of Transmission Eigenvalue Problem
In this section, we introduce the biharmonic scattering problem associated with the clamped transmission eigenvalues. Here, we assume that the clamped obstacle is modeled by a bounded region $D\subset\mathbb{R}^{2}$ with a Lipschitz boundary $\partial D$. To illuminate the obstacle, one uses a time–harmonic incident plane wave, which we denote as $u^{\text{inc}}(x)=\text{e}^{\text{i}kx\cdot d}$, where $d$ is the corresponding incident direction such that $d\in\mathbb{S}^{1}=\{x\in\mathbb{R}^{2}\mathrel{\mathop{\ordinarycolon}}|x|=1\}$. For our model, the corresponding scattered field $u^{\text{scat}}$ satisfies the biharmonic scattering problem for a fixed wave number $k>0$, given by
@@ 3 Existence of the Transmission Eigenvalues
We note that, for $k>0$ we have that both $u_{f}$ and $\partial_{r}u_{f}$ decay exponentially as $r\to\infty$ so no radiation condition at infinity is required but we assume that $u_{f}\in H^{1}(\mathbb{R}^{2}\setminus\overline{B_{R}})$. First, note that the eigenvalue problems (7)–(8) and (9)–(10) are equivalent.
@@
This implies that we can prove the existence of eigenpairs $k>0$ and $(v,w)\in X(D,B_{R})$ to (9)–(10) where the Hilbert space
@@
over the complex plane with the associated inner–product
@@
Therefore, to study the transmission eigenvalue problem in the truncated domain. By appealing to Green’s first identity we have that the equivalent variational formulation of the clamped transmission eigenvalue problem (9)–(10) is given by: find eigenpairs $k>0$ and $(v,w)\in X(D,B_{R})$ such that
@@
Notice that $k>0$ is a clamped transmission eigenvalue if and only if the operator $\mathbb{A}_{k}-k^{2}\mathbb{B}$ has a non–trivial null space.
As previously stated, in [24] it has been shown that the set of clamped transmission eigenvalues $k>0$ corresponding to (7)–(8) (or equivalently (9)–(10)) is at most discrete. With this, we focus on the existence of clamped transmission eigenvalues in this work. To prove this claim, we will study the analytic properties of the operators $\mathbb{A}_{k}$ and $\mathbb{B}$ defined in (16)–(17). Motivated by the results in [11], we will show that the aforementioned operators satisfy the assumptions of a key result. To this end, we recall Theorem 2.3 of [11] which will be applied to our eigenvalue problem.
@@
The operator $\mathbb{A}_{0}$ will be defined later in equation (19) by studying the limit as $k\to 0^{+}$ for the DtN mapping. This result is proven by considering the problem $\lambda_{j}(k)-k^{2}=0$ where $\lambda_{j}(k)$ are the generalized eigenvalues such that $\mathbb{A}_{k}-\lambda_{j}(k)\mathbb{B}$ has a non–trivial null space. Since each $\lambda_{j}(k)$ is continuous with respect to $k\geq 0$, the positivity and non–positivity requirements in Lemma 3.1 implies that
@@
Notice that $C_{k}$ can depend on $k>0$ but not $f$ and we make the dependance of the solution with respect to $k$ explicit. We will now employ a variational argument to prove that
@@
Now, to continue we let $u^{k}_{f}$ and $u^{\tau}_{f}$ satisfy the above boundary value problem with ‘wave numbers’ $k>0$ and $\tau>0$, respectively. Notice that
@@
Therefore, we obtain the estimate
@@
With the above analysis we can now prove that the operators $\mathbb{A}_{k}$ and $\mathbb{B}$ satisfy the assumptions of Lemma 3.1. We prove that each operator satisfies the assumptions in the following two results.
@@
$\left|\left(\left(\mathbb{A}_{{k}}-\mathbb{A}_{{\tau}}\right)(v,w);(\psi_{1},\psi_{2})\right)_{X(D,B_{R})}\right|=\left|\int_{B_{R}\setminus\overline{D}}\left(k^{2}-\tau^{2}\right)w\overline{\psi_{2}}\,\mathrm{d}x-\int_{\partial B_{R}}\overline{\psi_{2}}\big(T_{\text{i}k}-T_{\text{i}\tau}\big)w\,\mathrm{d}s\right|.$
@@
where $\|\cdot\|$ denotes the appropriate operator norms. By Theorem 3.2 we have that $\|\mathbb{A}_{{k}}-\mathbb{A}_{{\tau}}\|\longrightarrow 0$ as $\tau\to k$, proving the continuity.
Now, to prove that the operator is self–adjoint, by appealing to the definition of the sesquilinear form in (14) we notice that
$\left(\mathbb{A}_{k}(v,w);(v,w)\right)_{X(D,B_{R})}$ $=\int_{D}|\nabla v|^{2}\,\mathrm{d}x+\int_{B_{R}\setminus\overline{D}}{|\nabla w|^{2}+k^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}}}\,\mathrm{d}x-\int_{\partial B_{R}}\overline{w}T_{\text{i}k}w\,\mathrm{d}s$
$=\int_{D}|\nabla v|^{2}\,\mathrm{d}x+\int_{B_{R}\setminus\overline{D}}{|\nabla w|^{2}+k^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|w|^{2}}}\,\mathrm{d}x+\int_{\mathbb{R}^{2}\setminus\overline{D}}{|\nabla u_{w}|^{2}+k^{2}|u_{w}|^{2}}\mathrm{d}x.$
@@
Since the sequence is bounded, we have that (up to a subsequence) it is weakly convergent to some $(v,w)\in{X(D,B_{R})}$. This implies that $w_{n}\to w=0$ in $H^{1}(B_{R}\setminus\overline{D})$ and that $|\nabla v_{n}|\to|\nabla v|=0$ in $L^{2}(D)$ as $n\to\infty$. Therefore, we see that the limiting function $v$ is constant. By appealing to the compact embedding $H^{1}(D)$ into $L^{2}(D)$ we have that (up to a subsequence) $v_{n}$ is strongly convergent in $H^{1}(D)$. The boundary condition v=wv=-w on $\partial D$ implies that $v=0$ in $D$. The convergence $(v_{n},w_{n})\to(0,0)$ in $X(D,B_{R})$ as $n\to\infty$ contradicts the fact that $\|(v_{n},w_{n})\|_{X(D,B_{R})}=1$. This proves the coercivity and therefore proves the claim. ∎
@@
It is clear that from the sequilinear form (15) that $\mathbb{B}$ given by (17) is self–adjoint and non–negative since $\mathcal{B}\big((v,w)\,;(v,w)\big)=\|v\|^{2}_{L^{2}(D)}$. For the compactness, by definition we can easily obtain that $\|\mathbb{B}(v,w)\|_{X(D,B_{R})}\leq\|v\|_{L^{2}(D)}$. The compact embedding of $H^{1}(D)$ into $L^{2}(D)$ proves the claim, since $D$ is a bounded domain with a Lipschitz boundary. ∎
@@
where, $H^{(1)}_{\ell}$ denotes the first kind Hankel function of order $\ell$ and $f_{n}$ are the Fourier coefficients for $f\in H^{1/2}(\partial B_{R})$. Note that above we have used the fact that
@@ 4 Numerical Investigation
is the radiating fundamental solution for the Helmholtz equation in $\mathbb{R}^{2}$. Here, $\lambda$ and $\mu$ are the corresponding Dirichlet and Neumann eigenvalues when
@@ 5 Conclusion
In conclusion, we have provided an analytical and numerical study of the new clamped transmission eigenvalue problem. This eigenvalue problem is associated with scattering in a thin elastic plate. Even though the scatterer is bounded, the eigenvalue problem (7)–(8) is posed in the entirety of $\mathbb{R}^{2}$. With the results in this paper, we now know that there exists an infinite discrete set of real eigenvalues such that $k>0$. Here, we have provided more justification that the first clamped transmission eigenvalue is monotonically decreasing with respect to the measure of $D$, but it has not yet been proved. Another question that arises from Theorem 3.8 is if there are interlacing inequalities with other eigenvalues associated with the scatterer, which we have numerically investigated in Section 4. Also, the eigenvalue problem studied here is derived via a clamped obstacle assumption. It would be interesting to see what eigenvalue problems one derives from either the simply supported or free plate obstacle assumption. Lastly, spectral patterns of transmission eigenfunctions have received considerable attentions with striking applications in [2, 14, 15, 19] which can also be studied for this model.